OpenStudy (anonymous):

Can someone help me? I know the answer is A but can someone show me how? Thanks! Solve. 2/ X+3 < 0 A x < –3 B x > 3 C x > –3 D x < 3

OpenStudy (amistre64):

\[\frac 2x+3<0\] or \[\frac 2{x+3}<0\] ???

OpenStudy (anonymous):

2/ X+3 < 0

OpenStudy (amistre64):

that doesnt clear it up any for me

OpenStudy (anonymous):

no the second one!

OpenStudy (anonymous):

do u divide 2 by x+3

OpenStudy (amistre64):

lol, thats better :)

OpenStudy (anonymous):

haah sorry

OpenStudy (amistre64):

since there is not actual "0" value, we have to consider the values that make the bottom go zero

OpenStudy (amistre64):

since x+3 = 0 when x=-3 we need to check "sides" now

OpenStudy (amistre64):



OpenStudy (amistre64):

when x=0, 2/3 < 0 is false which means the other side is true by default